OpenStudy (anonymous):

show all your work explain how you got these answers 4t+5=37 r/2-8=-4 5m-8=57 show all work for each person please.

OpenStudy (anonymous):

1) 4t+5=37 4t=37-5 4t=32 divide both sides by 4. t=8. 2) r/2-8=-4 r/2=-4+8 r/2=4 multiply both sides by 2. r=8. 3) 5m-8=57 5m=57+8 5m=65 divide both sides by 5. m=13. :)